please reply with only the number answer to the underlined question. i do not want to see any other equations or numbers. just the answer to the underlined questions. thank you.
Question 3 1 pts In order for your clients to reach their retirement goal, lets say youve estimated that they need to have
Answer #1

Since you are only asking for the number:

The amount that needs to be saved at the end of each month is $4822.80

Just in case the detailed solution is required;
The real rate of return after adjusting for inflation is (1.079/1.028) -1 = 4.96%

So the monthly rate of return is 4.96/12 = 0.4133

Now they are currently 40 years old and plan to retire at 61, so number of months is (61-40) * 12 = 252

Present value (savings at present) 130,000

Interest rate = 0.4133

Future value of saving =2500000

we have calculate the monthly savings, (PMT), use financial calculator

PMT comes out to be$ 4822.79